# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ IV.6 Computing $R_{,v}$
$R_{,v}(u,v)=-\frac{\widetilde{T}_{vv}^{(-)}(v)}{\kappa_{-}^{v}(v)}$ (IV.40)
IV.7 Computing $R_{,u}$
We next focus on $R_{,u}$. We recall that $\widetilde{T}_{uu}\equiv 8\pi r^{2}T_{uu}=R_{,u}S_{,u}-R_{,uu}$, and we make a similar ansatz that $R_{,uu}\ll R_{,u}S_{,u}$ because $R_{,uu}\propto\varepsilon^{2}$ while $R_{,u}\propto\varepsilon$ (justifying it is similar to the $R_{,v}$ case discussed above). Therefore, using Eq. (IV.16), and recalling that in region 3, $\widetilde{T}_{uu}=\widetilde{T}_{uu}^{(-)}(v=-u)$, we get:
$R_{,u}(u,v)=-\frac{\widetilde{T}_{uu}^{(-)}(v=-u)}{\kappa_{-}^{u}(u)}$ (IV.41)
To avoid confusion we should note the following: $\widetilde{T}_{uu}^{(-)}$ was defined as a function of $v$ (as the near-IH $\widetilde{T}_{uu}$ value in RN with parameters $M(v)$ and $Q(v)$). We also saw that in region 3, $\widetilde{T}_{uu}$ depends only on $u$ (see subsection IV.5). The expression we used above $\widetilde{T}_{uu}=\widetilde{T}_{uu}^{(-)}(v=-u)$ is indeed consistent with both of these requirements.
IV.8 Summary
Eqs. (IV.5,IV.16,IV.40,IV.41) describe the analytical approximation for the quantities $S_{,v}$, $S_{,u}$, $R_{,v}$, $R_{,u}$ in region 3. As can be seen in these equations, this analytical approximation only depends on a set of four semiclassical input functions: $M(v)$ and $Q^{v}(v)$ (or alternatively $\widetilde{T}_{vv}^{(+)}(v)$ and $J_{v}^{(+)}(v)$) at the event horizon, and $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ at the inner horizon. (Recall, these two last quantities are defined, via the semiclassical theory, as the near-IH fluxes in RN with parameters $M(v)$ and $Q(v)$, and are therefore viewed as functions of $v$.)
V Numerical algorithm
We would like to integrate the evolution equations Eq. (II.38,II.39) numerically. We do that by putting both $R$ and $S$ on a discrete grid. We denote the distance between grid points in the $u$ direction as $du$ and the distance between grid points in the $v$ direction as $dv$. The grid points are accessed by two integer indices $i,j$. For simplicity we chose $du=dv$ throughout this work.
The basic numerical algorithm computes $R$ at the point $(i,j)$ (marked red in Figure V.1) as a combination of 3 previous points (marked blue), and of the derivative $R_{,uv}$ evaluated at the middle point $(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (marked green), as illustrated in Figure V.1, as follows:
Figure V.1: Points in numerical algorithm
$R(i,j)=R(i-1,j)+R(i,j-1)-R(i-1,j-1)+du\;dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (V.1)
@@ V Numerical algorithm
Such a discrete numerical algorithm necessarily introduces a numerical error, called truncation error, which in this specific algorithm should be proportional to $du\cdot dv=du^{2}$. We can measure this numerical error by running the algorithm with two different values of $du$ and comparing the results obtained in these two runs, see subsection IX.1.
VI Initial values
In the vicinity of the EH, we presumably have, to a good approximation, a charged Vaidya solution, which evaporates at the rate of $\varepsilon$. Recall Eq. (II.100):
$ds^{2}=-\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}$ (VI.1)
In order to connect to the formalism developed above in subsection II.1 we have to transform it into double null coordinates. Unfortunately, we cannot implement this transformation analytically. Nevertheless, since $\varepsilon\ll 1$, at each point in this Vaidya geometry there is a local approximately-RN solution. Our strategy will be to use this fact to construct the (approximate) metric for the Vaidya geometry in double null coordinates.
VI.1 “naive $g_{1}$ metric”
We choose a certain $v$ value which we denote $v_{0}$. Due to the smallness of $\varepsilon$ there is a range of $v$ in the vicinity of $v_{0}$ in which the geometry is RN to a good approximation. The Vaidya metric there reduces to the RN geometry in ingoing Eddington coordinates $v,r$. We now transform it to double null coordinates $u,v$. Taking Eq. (II.80) and applying with local $M(v_{0})$ and $Q(v_{0})$ parameters:
$ds^{2}=\left(1-\frac{2M(v_{0})}{r(u,v)}+\frac{Q(v_{0})^{2}}{r(u,v)^{2}}\right)du\;dv+r(u,v)^{2}d\Omega^{2}$ (VI.2)
@@ VI Initial values
$r_{*}=r-M(v_{0})+\frac{1}{2\kappa_{+}(v_{0})}\log\left|\frac{r-r_{+}(v_{0})}{M(v_{0})-r_{+}(v_{0})}\right|-\frac{1}{2\kappa_{-}(v_{0})}\log\left|\frac{r-r_{-}(v_{0})}{M(v_{0})-r_{-}(v_{0})}\right|$ (VI.3)
@@
$ds^{2}=\left(1-\frac{2M(v)}{r(u,v)}+\frac{Q(v)^{2}}{r(u,v)^{2}}\right)du\;dv+r(u,v)^{2}d\Omega^{2}$ (VI.4)
@@
$\frac{u+v}{2}\equiv r_{*}=r-M(v)+\frac{1}{2\kappa_{+}(v)}\log\left|\frac{r-r_{+}(v)}{M(v)-r_{+}(v)}\right|-\frac{1}{2\kappa_{-}(v)}\log\left|\frac{r-r_{-}(v)}{M(v)-r_{-}(v)}\right|$ (VI.5)
@@
$\left(\frac{\partial r}{\partial r_{*}}\right)_{v}=f$ (VI.7)
The metric in Eq. (VI.4) provides a description of the Vaidya metric Eq. (VI.1) in double null coordinates, up to deviation of order $\varepsilon\ll 1$. We shall now discuss the regularity of this metric at the EH.
@@
$U=e^{\kappa_{+}u},\hfill V=e^{\kappa_{+}v}$ (VI.8)
Attempting to generalize these Kruskal coordinates to our case, since $\kappa_{+}$ is a function of $v$ (unlike the RN case) $U=e^{\kappa_{+}(v)u}$ is not a function of only $u$, and therefore it is not a null coordinate. Even worse, it turns out that the naive metric (VI.4) geometry is not regular at the EH: we will show in Appendix A that the Ricci scalar actually diverges there. To solve this problem we will move to the metric $\widetilde{g_{1}}$ (see next subsection).
Note that this problem arose when we moved from $v_{0}$ in the local RN metric Eq. (VI.2) to the global $v$-dependent metric (VI.4). We need to implement this transition in a different manner to avoid this pitfall, as we will now describe.
VI.2 The $\widetilde{g_{1}}$ metric
First, we will take the local Reissner–Nordström metric (at $v=v_{0}$) in double null coordinates Eq. (VI.2), and transform it to different double null coordinates $\widetilde{u}(u),\widetilde{v}(v)$ defined by:
$\widetilde{v}\equiv\stackrel{{\scriptstyle[}}{{0}}]{v}{\int}\kappa_{+}(v_{0})dv,\qquad\widetilde{u}\equiv\stackrel{{\scriptstyle[}}{{0}}]{u}{\int}\kappa_{+}(v_{0})du$ (VI.9)
@@ VI.2 The $\widetilde{g_{1}}$ metric
$ds^{2}=\frac{1}{\kappa_{+}(v_{0})^{2}}\left(1-\frac{2M(v_{0})}{r}+\frac{Q(v_{0})^{2}}{r^{2}}\right)d\widetilde{u}\;d\widetilde{v}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (VI.10)
@@
$\widetilde{r_{*}}=\frac{\widetilde{v}+\widetilde{u}}{2}$ (VI.11)
@@
$ds^{2}=\frac{1}{\kappa_{+}(\widetilde{v})^{2}}\left(1-\frac{2M(\widetilde{v})}{r}+\frac{Q(\widetilde{v})^{2}}{r^{2}}\right)d\widetilde{u}\;d\widetilde{v}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (VI.13)
@@
$\frac{\widetilde{u}+\widetilde{v}}{2}\equiv\widetilde{r_{*}}=\kappa_{+}(v)\left(r-M(v)+\frac{1}{2\kappa_{+}(v)}\log\left|\frac{r-r_{+}(v)}{M(v)-r_{+}(v)}\right|-\frac{1}{2\kappa_{-}(v)}\log\left|\frac{r-r_{-}(v)}{M(v)-r_{-}(v)}\right|\right)$ (VI.14)
@@
Next we discuss the regularity of this geometry at the EH. This metric, as is, still isn’t regular, because $g_{\widetilde{u}\widetilde{v}}=0$ at the EH. However, like in RN, we can transform it to Kruskal coordinates (identical to those defined in Eq. (VI.8)) which now take the form:
$U=e^{\widetilde{u}},\qquad V=e^{\widetilde{v}}$ (VI.15)
Recalling that our metric is just like RN along a line of constant $v$, and recalling Eq. (II.99) , the metric function $g_{\widetilde{u}\widetilde{v}}$ around $r_{+}$ will be
$g_{\widetilde{u}\widetilde{v}}=\frac{g_{uv}}{\kappa_{+}^{2}(v)}=-const(v)\cdot e^{\kappa_{+}\left(u+v\right)}=-const(v)\cdot e^{\widetilde{u}+\widetilde{v}}$ (VI.16)
where $const(v)$ is a non-zero constant which depends only on $v$. Transforming into $U,V$ we get:
$g_{UV}=-const(v)$ (VI.17)
@@
$r_{*}\equiv\frac{\widetilde{r_{*}}}{\kappa_{+}^{v}(v)}=\frac{\widetilde{u}+\widetilde{v}}{2\kappa_{+}^{v}(v)}$ (VI.19)
Therefore, using Eq. (VI.7), the derivative of $r$ with respect to $\widetilde{u}$ is:
$\frac{\partial r}{\partial\widetilde{u}}=\left(\frac{\partial r}{\partial r_{*}}\right)_{v}\left(\frac{\partial r_{*}}{\partial\widetilde{u}}\right)_{v}=\frac{f}{2\kappa_{+}^{v}(v)}$ (VI.20)
@@
$R_{,\widetilde{u}}=2rr_{,\widetilde{u}}=\frac{rf}{\kappa_{+}(v)}$ (VI.21)
@@
$\frac{e^{\widetilde{S}}}{r}=-g_{\widetilde{u}\widetilde{v}}=-\frac{f}{2\kappa_{+}(v)^{2}}$ (VI.22)
@@
$\widetilde{S}=\log\left(-\frac{rf}{2\kappa_{+}^{2}(v)}\right)$ (VI.23)
@@
$\widetilde{S}_{,\widetilde{u}}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)r_{,\widetilde{u}}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)\frac{f}{2\kappa_{+}^{2}}$ (VI.24)
@@
VI.3 $g_{1}$ metric
We wish to do coordinate transformation from $\widetilde{g_{1}}$ with $\widetilde{v},\widetilde{u}$ coordinates to new double null coordinates whose scaling (with regards to $M$) is more similar to the Eddington $v,u$ coordinates in RN. (Because it is just coordinate transformation, it will keep the geometry regular at the EH.) For simplicity, we’ll call these new coordinates $v,u$ (however, this new metric should not be confused with the “naive metric” described in subsection VI.1). We shall call the new resulting metric simply “the $g_{1}$ metric”.

### v2

@@ 4.6 Computing $R_{,v}$
$R_{,v}(u,v)=-\frac{\widetilde{T}_{vv}^{(-)}(v)}{\kappa_{-}^{v}(v)}$ (4.40)
4.7 Computing $R_{,u}$
We next focus on $R_{,u}$. We recall that $\widetilde{T}_{uu}\equiv 8\pi r^{2}T_{uu}=R_{,u}S_{,u}-R_{,uu}$, and we make a similar ansatz that $R_{,uu}\ll R_{,u}S_{,u}$ because $R_{,uu}\propto\varepsilon^{2}$ while $R_{,u}\propto\varepsilon$ (justifying it is similar to the $R_{,v}$ case discussed above). Therefore, using Eq. (4.16), and recalling that in region 3, $\widetilde{T}_{uu}=\widetilde{T}_{uu}^{(-)}(v=-u)$, we get:
$R_{,u}(u,v)=-\frac{\widetilde{T}_{uu}^{(-)}(v=-u)}{\kappa_{-}^{u}(u)}$ (4.41)
To avoid confusion we should note the following: $\widetilde{T}_{uu}^{(-)}$ was defined as a function of $v$ (as the near-IH $\widetilde{T}_{uu}$ value in RN with parameters $M(v)$ and $Q(v)$). We also saw that in region 3, $\widetilde{T}_{uu}$ depends only on $u$ (see subsection 4.5). The expression we used above $\widetilde{T}_{uu}=\widetilde{T}_{uu}^{(-)}(v=-u)$ is indeed consistent with both of these requirements.
4.8 Summary
Eqs. (4.5,4.16,4.40,4.41) describe the analytical approximation for the quantities $S_{,v}$, $S_{,u}$, $R_{,v}$, $R_{,u}$ in region 3. As can be seen in these equations, this analytical approximation only depends on a set of four semiclassical input functions: $M(v)$ and $Q^{v}(v)$ (or alternatively $\widetilde{T}_{vv}^{(+)}(v)$ and $J_{v}^{(+)}(v)$) at the event horizon, and $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ at the inner horizon. (Recall, these two last quantities are defined, via the semiclassical theory, as the near-IH fluxes in RN with parameters $M(v)$ and $Q(v)$, and are therefore viewed as functions of $v$.)
5 Numerical algorithm
We would like to integrate the evolution equations Eq. (2.37,2.38) numerically. We do that by putting both $R$ and $S$ on a discrete grid. We denote the distance between grid points in the $u$ direction as $du$ and the distance between grid points in the $v$ direction as $dv$. The grid points are accessed by two integer indices $i,j$. For simplicity we chose $du=dv$ throughout this work.
The basic numerical algorithm computes $R$ at the point $(i,j)$ (marked red in Figure 5.1) as a combination of 3 previous points (marked blue), and of the derivative $R_{,uv}$ evaluated at the middle point $(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (marked green), as illustrated in Figure 5.1, as follows:
Figure 5.1: Points in numerical algorithm
$R(i,j)=R(i-1,j)+R(i,j-1)-R(i-1,j-1)+du\;dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (5.1)
@@ 5 Numerical algorithm
Such a discrete numerical algorithm necessarily introduces a numerical error, called truncation error, which in this specific algorithm should be proportional to $du\cdot dv=du^{2}$. We can measure this numerical error by running the algorithm with two different values of $du$ and comparing the results obtained in these two runs, see subsection 9.1.
6 Initial values
In the vicinity of the EH, we presumably have, to a good approximation, a charged Vaidya solution, which evaporates at the rate of $\varepsilon$. Recall Eq. (2.99):
$ds^{2}=-\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}$ (6.1)
In order to connect to the formalism developed above in subsection 2.1 we have to transform it into double null coordinates. Unfortunately, we cannot implement this transformation analytically. Nevertheless, since $\varepsilon\ll 1$, at each point in this Vaidya geometry there is a local approximately-RN solution. Our strategy will be to use this fact to construct the (approximate) metric for the Vaidya geometry in double null coordinates.
6.1 “naive $g_{1}$ metric”
We choose a certain $v$ value which we denote $v_{0}$. Due to the smallness of $\varepsilon$ there is a range of $v$ in the vicinity of $v_{0}$ in which the geometry is RN to a good approximation. The Vaidya metric there reduces to the RN geometry in ingoing Eddington coordinates $v,r$. We now transform it to double null coordinates $u,v$. Taking Eq. (2.79) and applying with local $M(v_{0})$ and $Q(v_{0})$ parameters:
$ds^{2}=\left(1-\frac{2M(v_{0})}{r(u,v)}+\frac{Q(v_{0})^{2}}{r(u,v)^{2}}\right)du\;dv+r(u,v)^{2}d\Omega^{2}$ (6.2)
@@ 6 Initial values
$r_{*}=r-M(v_{0})+\frac{1}{2\kappa_{+}(v_{0})}\log\left|\frac{r-r_{+}(v_{0})}{M(v_{0})-r_{+}(v_{0})}\right|-\frac{1}{2\kappa_{-}(v_{0})}\log\left|\frac{r-r_{-}(v_{0})}{M(v_{0})-r_{-}(v_{0})}\right|$ (6.3)
@@
$ds^{2}=\left(1-\frac{2M(v)}{r(u,v)}+\frac{Q(v)^{2}}{r(u,v)^{2}}\right)du\;dv+r(u,v)^{2}d\Omega^{2}$ (6.4)
@@
$\frac{u+v}{2}\equiv r_{*}=r-M(v)+\frac{1}{2\kappa_{+}(v)}\log\left|\frac{r-r_{+}(v)}{M(v)-r_{+}(v)}\right|-\frac{1}{2\kappa_{-}(v)}\log\left|\frac{r-r_{-}(v)}{M(v)-r_{-}(v)}\right|$ (6.5)
@@
$\left(\frac{\partial r}{\partial r_{*}}\right)_{v}=f$ (6.7)
The metric in Eq. (6.4) provides a description of the Vaidya metric Eq. (6.1) in double null coordinates, up to deviation of order $\varepsilon\ll 1$. We shall now discuss the regularity of this metric at the EH.
@@
$U=e^{\kappa_{+}u},\hfill V=e^{\kappa_{+}v}$ (6.8)
Attempting to generalize these Kruskal coordinates to our case, since $\kappa_{+}$ is a function of $v$ (unlike the RN case) $U=e^{\kappa_{+}(v)u}$ is not a function of only $u$, and therefore it is not a null coordinate. Even worse, it turns out that the naive metric (6.4) geometry is not regular at the EH: we will show in Appendix A that the Ricci scalar actually diverges there. To solve this problem we will move to the metric $\widetilde{g_{1}}$ (see next subsection).
Note that this problem arose when we moved from $v_{0}$ in the local RN metric Eq. (6.2) to the global $v$-dependent metric (6.4). We need to implement this transition in a different manner to avoid this pitfall, as we will now describe.
6.2 The $\widetilde{g_{1}}$ metric
First, we will take the local Reissner–Nordström metric (at $v=v_{0}$) in double null coordinates Eq. (6.2), and transform it to different double null coordinates $\widetilde{u}(u),\widetilde{v}(v)$ defined by:
$\widetilde{v}\equiv\mathrel{{\mathop{\int}\limits_{0}^{v}}}\kappa_{+}(v_{0})dv,\qquad\widetilde{u}\equiv\mathrel{{\mathop{\int}\limits_{0}^{u}}}\kappa_{+}(v_{0})du$ (6.9)
@@ 6.2 The $\widetilde{g_{1}}$ metric
$ds^{2}=\frac{1}{\kappa_{+}(v_{0})^{2}}\left(1-\frac{2M(v_{0})}{r}+\frac{Q(v_{0})^{2}}{r^{2}}\right)d\widetilde{u}\;d\widetilde{v}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (6.10)
@@
$\widetilde{r_{*}}=\frac{\widetilde{v}+\widetilde{u}}{2}$ (6.11)
@@
$ds^{2}=\frac{1}{\kappa_{+}(\widetilde{v})^{2}}\left(1-\frac{2M(\widetilde{v})}{r}+\frac{Q(\widetilde{v})^{2}}{r^{2}}\right)d\widetilde{u}\;d\widetilde{v}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (6.13)
@@
$\frac{\widetilde{u}+\widetilde{v}}{2}\equiv\widetilde{r_{*}}=\kappa_{+}(v)\left(r-M(v)+\frac{1}{2\kappa_{+}(v)}\log\left|\frac{r-r_{+}(v)}{M(v)-r_{+}(v)}\right|-\frac{1}{2\kappa_{-}(v)}\log\left|\frac{r-r_{-}(v)}{M(v)-r_{-}(v)}\right|\right)$ (6.14)
@@
Next we discuss the regularity of this geometry at the EH. This metric, as is, still isn’t regular, because $g_{\widetilde{u}\widetilde{v}}=0$ at the EH. However, like in RN, we can transform it to Kruskal coordinates (identical to those defined in Eq. (6.8)) which now take the form:
$U=e^{\widetilde{u}},\qquad V=e^{\widetilde{v}}$ (6.15)
Recalling that our metric is just like RN along a line of constant $v$, and recalling Eq. (2.98) , the metric function $g_{\widetilde{u}\widetilde{v}}$ around $r_{+}$ will be
$g_{\widetilde{u}\widetilde{v}}=\frac{g_{uv}}{\kappa_{+}^{2}(v)}=-W(v)\cdot e^{\kappa_{+}\left(u+v\right)}=-W(v)\cdot e^{\widetilde{u}+\widetilde{v}}$ (6.16)
where $W(v)$ is a non-zero function which depends only on $v$. Transforming into $U,V$ we get:
$g_{UV}=-W(v)$ (6.17)
@@
$r_{*}\equiv\frac{\widetilde{r_{*}}}{\kappa_{+}^{v}(v)}=\frac{\widetilde{u}+\widetilde{v}}{2\kappa_{+}^{v}(v)}$ (6.19)
Therefore, using Eq. (6.7), the derivative of $r$ with respect to $\widetilde{u}$ is:
$\frac{\partial r}{\partial\widetilde{u}}=\left(\frac{\partial r}{\partial r_{*}}\right)_{v}\left(\frac{\partial r_{*}}{\partial\widetilde{u}}\right)_{v}=\frac{f}{2\kappa_{+}^{v}(v)}$ (6.20)
@@
$R_{,\widetilde{u}}=2rr_{,\widetilde{u}}=\frac{rf}{\kappa_{+}(v)}$ (6.21)
@@
$\frac{e^{\widetilde{S}}}{r}=-g_{\widetilde{u}\widetilde{v}}=-\frac{f}{2\kappa_{+}(v)^{2}}$ (6.22)
@@
$\widetilde{S}=\log\left(-\frac{rf}{2\kappa_{+}^{2}(v)}\right)$ (6.23)
@@
$\widetilde{S}_{,\widetilde{u}}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)r_{,\widetilde{u}}=\left(\frac{\partial}{\partial r}\log\left(-rf\right)\right)\frac{f}{2\kappa_{+}^{2}}$ (6.24)
@@
6.3 $g_{1}$ metric
We wish to do coordinate transformation from $\widetilde{g_{1}}$ with $\widetilde{v},\widetilde{u}$ coordinates to new double null coordinates whose scaling (with regards to $M$) is more similar to the Eddington $v,u$ coordinates in RN. (Because it is just coordinate transformation, it will keep the geometry regular at the EH.) For simplicity, we’ll call these new coordinates $v,u$ (however, this new metric should not be confused with the “naive metric” described in subsection 6.1). We shall call the new resulting metric simply “the $g_{1}$ metric”.